### v1

@@ Abstract
It is worth noting that a hole in the ligand orbital is a source of instability which a positive charge transfer system often corrects, annihilating the hole, in time intervals of femtosecond via radiation or thermal vibration. However, in negative charge transfer materials, the electronic structure modifications near manifest as a correlated redistribution of charge, leading to various newly-formed hybridised states where the hole coexists with the compensating Ni 3d charge. We now have the tools to investigate redox in LiNiO2.
@@
The analysis in previous section provides the analytical tools for an in-depth investigation of the redox process in LiNiO2 cathodes. We prepared half coin cells (vs. Li/Li+) and charged them at C/20 to three different upper cut-off voltages (4.2, 4.4 and 4.6 V), as shown by the electrochemical profiles in Figure 3b. Similar to the O K-edge spectra for the set of Ni oxides in Figure 2d for the negative materials, increasing the Ni oxidation state in the LNO cathode as a result of delithiation correlates, in Figure 3a, with an increasing O K-edge pre-edge peak, which the previous section demonstrates to be sensitive to the ligand hole formation. The Ni L-edge assisted RPES demonstrates the nature of this orbital and how it evolves with the cell potential. Figure 3c shows the Ni L-edge experiments and the energy region (855 - 865 eV) of the Ni L-edge transitions related to the strong hybridised Ni 3d and O ligand 2p orbital with stable ligand holes. Resonating within this region in RPES studies enhances orbitals with ligand hole character which appears at approximately 2 eV. These orbitals’ intensity, located close to the level, reveals a mild withdrawal of electrons (white square in Figure 3d) which evolves with higher cell potentials, indicating charge compensation occurs in these ligand hole containing orbitals. In contrast, RPES shows an enhanced Ni 3d orbital signal at approximately 10 eV that doesn’t evolve with the electrochemistry. Taking together, oxidation state evolution of LNO electrodes does not translate to valence-counting of the Ni 3d orbital but rather is tightly linked to the evolution of the ligand hole containing Ni 3d - O ligand 2p hybridised orbital. Elucidating the exact nature of this hybridised orbital and a clear redox picture within the ligand hole picture is achieved using SIA simulations of the Ni L-edge XAS experiments.
Determining the correct ligand hole configuration at the beginning of the charge process is critical to elucidating the redox mechanism. We start by considering as the main ligand hole configuration in pristine LiNiO2. When one reads this configuration in the literature, a couple of conditions are often overlooked. First, the spatial distribution of this hybridised orbital is assumed to be confined to a single NiO6 site. And second, due to this single-site spatial extension, the state results from the hybridisation between the Ni 3d eg orbitals and the O ligand orbitals that directly point to the Ni ion as seen in Figure 4a. Taking together, these two conditions lead to a NiO6 layer configuration where every single octahedra adopts the ligand configuration as seen in Figure 4b. The overlap between the theory with these conditions and the experimental data is suboptimal in Figure 4c because it is limited to the isolated NiO6 unit and does not consider the entire crystallographic structure.
@@
A more realistic approach is considering that the orbital hybridisation expands across the entire structure as in the case of band theory. Indeed, when the Ni -O ligand hybridisation extends across multiple NiO6 octahedra, it leads to a more efficient way for the material to stabilise the ligand hole via a “self-regulating response”. [40] This indicates that transferring ligand electrons to the metal is regulated collectively within a set of neighbouring NiO6 octahedra in which the ligand holes reorganise among them via ligand-metal bond alternation, leading to configurations of the type seen in Figure 4e. These configurations have been observed in structural experiments of LiNiO2. [18, 20] As a result, we have ligand configurations of the type , , and that correspond geometrically to unique octahedral configurations. Our hypothesis of the LiNiO2 electronic configuration having an equal distribution of these three types replicates well the XAS spectra, as shown in Figure 4f, confirming the self-regulating response is a realistic electronic structure for the pristine LiNiO2 electrode. We simulated Ni L-edge transitions with the three electronic configurations in three independent SIA models. And although the highly covalent yet correlated nature of LiNiO2 requires a more complex single Ni L-edge simulation with multiple clusters of NiO6 to properly capture bond-alternation, correlation, and covalency at once, [41, 42] the individual SIA approach herein adopted provides key insights into the basic nature of the Ni L-edge transitions and the corresponding experimental peaks.
@@
Understanding the redox mechanism at the LiNiO2 cathode side during the charging process ultimately comes down to clarifying how the three , , and ligand configurations adjust during intercalation reactions. In this sense, we have run bulk sensitive Ni L-edge XAS measurements at different voltage points along the electrochemical profile of half coin cells (also charged at a C-rate of C/20), which are highlighted in Figure 5a. From our previous calculations in Figure 4f we found that the XAS spectra can be interpreted in terms of the intensity of three peaks (at roughly 853, 854 and 855 eV), which denote the distribution within the LiNiO2 crystallographic structure of three different types of NiO6 sites and ligand hole configurations. As the battery is charged to higher voltages, up to 4.4 V, there is a gradual increase in the intensity of the peak, while the peak is reduced. This suggests that the evolution of the electronic configuration during charge is characterised by an increased density of ligand holes at the oxygen sites, in alignment with the XAS data at the O K-edge in Figure 3a. While the cell charging process relates to ligand hole creation and an electronic redistribution to allocate those additional ligand holes, the filling of these ligand holes dominates the discharge process which we can see in the spectral reduction of the and the increasing spectral features of the configuration after discharging from 4.4 V to 3.0 V vs. Li+/Li. A direct consequence of this creation/filling dynamics is the response of the local geometries adapting to the hole evolution. A follow-up study will investigate this phenomenon more in-depth using operando Ni K-edge XAS to elucidate the local range evolution in the context of redox evolution with ligand holes. Instead of counting Ni d electrons, in negative charge transfer compounds, the redox mechanism refers to counting those hole-induced distorted octahedral sites, caused by the dynamical evolution of ligand electronic density.
@@
The precise description of the electronic structure of LMFP64 and LNO is crucial for developing an understanding of the corresponding redox mechanism and oxidation evolution during intercalation reactions of the materials. It might have seemed that we defined the LMFP64 and LNO electronic structure in our SIA TM L-edge XAS simulations guided solely on simulations that match experimental measurements. However, we used the Zaan-Sawatzky-Allen (ZSA) electronic scheme [43] to infer the LMFP64 and LNO electronic structures. The ZSA framework classifies materials into four categories based on the material’s relative energy values between and U, which is the on-site Coulomb interaction and is the electron repulsion causing a separation in energy of the 3d-derived electronic states.
While repulsive Coulomb effects tend to localised charge around the metal centre, charge transfer favours delocalisation. This competing effect defines the ZSA category of the material, the resulting character of the charge-compensating orbitals, and consequently the charge compensation mechanism. To exemplify this, let’s look at the material band gap type formed within the ZSA electronic framework for LMFP64 and LNO, illustrating how this band gap can aid in determining the expected material redox nature (Figure 6).
@@
Model A in Figure 6 describes the ZSA Mott-Hubbard material class—positive ( U)—applicable to the LMFP64 case. The band gap results from Coulomb repulsion, so it is of a d - d nature. This means that the lowest unoccupied band (conduction band) and highest occupied band (valence band) are both of Fe d character. During cell charge, charge from the Fe 3d valence band will be promoted to the conduction band. The loss of Fe intensity in Figure 1d is a manifestation of this effect - hence Fe oxidation. Once in the conduction band, this promoted charge can reach the current collector of the cathode, leading to the measured current during the cell charge process. Model B in Figure 6 defines the ZSA Negative Charge Transfer material class which determines the LNO electronic structure. Since in these materials, is smaller than U and negative, the band gap is of a p - p type. During the charge process of LNO cells, electrons are promoted between this p - p band gap, creating holes in the ligand orbitals while drawing capacity. This system instantaneously re-structures the bands, sensitive to Ni-O bonds resulting in the reconfiguration of the distribution of the , , and states during electrochemistry. As a result, the occupancy of the Ni 3d band does not change, explaining the invariant intensity of the Ni L-edge assisted RPES heatmaps.
@@
This study encourages extending this framework to lithium excess compounds—the original systems that spurred interest in oxygen redox mechanisms to begin with. Note that the transition from positive to negative charge transfer regimes depends on metal and its formal oxidation state, corresponding to roughly Ni2+, Co3+ and Mn4+, [43] and that the majority of Li-excess compounds are Li-rich Mn-rich. [3] Thus, we anticipate these materials could be classed in one of the other two ZSA categories (Positive Charge Transfer or Mixed Valence) with varying p-d band gap types, [16] offering a pathway to study the role of unhybridised oxygen orbitals in delivering reversible capacity without conflicting signals. Li2MnO3 and Li2NiO3 reflect the metal and ligand hole endpoints, respectively, consistent with the observed electrochemistry arising from structural and chemical decomposition. [44, 37] We acknowledge that structural rearrangements also play a key role in the electrochemical behaviour of Li-rich systems. [22] TM L-edge assisted RPES offers an opportunity to directly explore electronic structures of Li-excess compounds under the ZSA scheme during (de)lithiation and to validate models and descriptions that clarify the role of oxygen oxidation in these compounds.
@@ Conclusion
Here, we present how by experimentally capturing the true nature of the electronic structure of LiMn0.6Fe0.4PO4 and LiNiO2 cathode materials under the ZSA scheme, the TM L-edge assisted RPES approach can directly probe their corresponding redox mechanism. Using this framework, we can conclude that the degree of metal-oxygen hybridisation is a key factor in determining the charge compensation mechanism. For more ionic compounds in the positive charge regime, such as the phosphates, electron counting of the d orbital remains valid (e.g. Fe2+/3+ 3d6,3d5). Increasingly covalent systems in the negative charge transfer regime, which includes the LiMO2 (M= Co, Ni) derived family, are better described in terms of electron counting ligand hole states instead (e.g. Ni3+/4+ , ). We expect this framework to also help describe lithium excess compounds. A more accurate probe into the evolution of their electronic structure may resolve the true role of unhybridised oxygen orbitals in delivering reversible capacity in these systems.
Supplementary information Not applicable
@@ Acknowledgements
We acknowledge Diamond Light Source for time on beamline B07 under proposal S33459-1 and beamline I09 under proposals SI35075-1, SI36917-1, and SI30201. This work was supported by the Faraday Institution’s Degradation (FIRG060, FIRG082) project.
Author Contributions G.J.P.F.: Conceptualization, Project Administration, Supervision, and Writing - Original Draft. D.D., M.A., and V.M.: Data curation, Methodology, Visualization, and Investigation. D.D.: Writing - Original Draft, Formal Analysis, and Validation. M.J.W.O., and M.A.: Writing - Review & Editing. M.H., S.H., P.P., M.I., B.K.: Resources - LaNiO2 single crystal synthesis and material quality control. I.M., and S.A.C.: Resources - LiNiO2 material synthesis and material quality control. P.K.T, T.L., D.C.G., P.F.: Investigation and Instrumental Resources - guidance in data collection during beamtimes. L.F.J.P.: Funding acquisition, Conceptualization, Project Administration, Supervision, Writing - Review & Editing
@@
The assembly of 2325 coin cells (vs. Li/Li+ as a counter electrode) was conducted in the glovebox. A Celgard 2325 Trilayer microporous membrane of 25 m thickness was used as a separator. The cells were filled with 60 L of E151 Solvionic electrolyte, comprised of 12.42:30.82:54.76:2 w/w ratio of LiPF6, ethylene carbonate (EC), ethyl methyl carbonate (EMC), and vinylene carbonate (VC), respectively. To ensure full wetting of the electrodes, the cells were held at rest at 25 ∘C) for 20 hours. Cells underwent one constant-current C/20 charge process to voltages (4.0 V, 4.1 V, 4.2 V, 4.3 V, 4.4 V, and 4.6 V) on a Biologic VMP3 potentiostat cycler.
@@
After required cycling coin cells were carefully disassembled within a glovebox to prevent exposure to atmospheric moisture. The electrodes were thoroughly washed using dimethyl carbonate (DMC) solvent (sourced from Sigma-Aldrich) with an anhydrous purity of 99 percent. After washing and drying, the cleaned electrodes were adhere on copper plates before securely transferred to synchrotron facility for detailed characterization. To ensure the protection from environmental contamination, they were sealed in a sample cassette, provided by diamond light source, UK.
@@
LiMn0.6Fe0.4PO4 (LMFP) powder was purchased from Gelon LIB and used to fabricate electrodes following in-house procedures. A mixture of LMFP powder, carbon black, and polyvinylidene fluoride (PVDF) pre-dissolved in N-methyl-2-pyrrolidone (NMP) was prepared in a weight ratio of 93.5:3.5:3.5. Additional NMP was incorporated to achieve the desired solid content. The resulting slurry was then applied to carbon-coated aluminium foil, dried overnight at 120 ∘C, and calendared to a final density of 2.15 g cm-3. To charge the material to a specific state of charge (SoC), half coin cells were assembled with LMFP cathodes (14.8 mm diameter), a polymer separator (H2325), LP57 electrolyte, and a lithium metal disc as the counter electrode. The cells were then charged galvanostatically to a specified potential using a current of 0.174 mA (equivalent to a C/20 charge rate). After charging, the cells were disassembled in an Ar-filled glovebox, and the electrodes were rinsed with DMC and dried. The electrodes were subsequently mounted on a sample holder and shipped under vacuum to the Diamond Light Source synchrotron.
@@
The valence-band data at Ni 2p absorption thresholds were recorded at beamline I09 Diamond Light Source (DLS) under the SI30201 beamtime with a VG Scienta EW4000 detector with a 70 frame/sec CCD camera. The total-energy resolution for the measurements at I09 was 0.2 eV. The calibration of the photon energy was performed by comparing the kinetic energy of the Au 4f peak. The base pressure during all measurements was less than 210-10 torr and all measurements were performed at room temperature.
@@
Ni L-edge X-ray Absorption Spectroscopy were measured in TEY and total fluorescent yield (TFY) mode at the B07 beamline [49] at DLS under the SI33459 beamtime. Coin cells held at different voltages were opened and the cathode electrode retrieved in inert argon conditions to avoid air exposure of the cathode surface. Fe L-edge X-ray Absorption Spectroscopy data were collected at the I09 beamline at DLS. In both cases, samples were transported to the beamline facilities and secured directly to the instrument antechamber with a sample transfer provided by the beamline scientists guaranteeing no-air exposure during the operation and transport of the material. The Ni and Fe L-edge spectra was collected with a energy resolution 0.15 and 0.2 eV respectively and energy scale was calibrated using Ni oxide and Iron oxide references.

### v3

@@ Abstract
It is worth noting that a hole in the ligand orbital is a source of instability which a positive charge transfer system often corrects, annihilating the hole, in time intervals of femtosecond via radiation or thermal vibration. However, in negative charge transfer materials, the electronic structure modifications near manifest as a correlated redistribution of charge, leading to various newly-formed hybridised states where the hole coexists with the compensating Ni 3d charge.
@@
The analysis presented in previous section provides a robust framework to investigate the redox process in LNO positive electrodes. In Fig. S4, as the cell potential increases, the O K-edge pre-edge feature, previously associated with ligand hole formation, exhibits a concurrent intensity growth (Fig. S4b). This trend suggests that since the pristine material already exhibits this pre-edge spectral feature, the electrochemical charging process, across the full voltage range, is governed by the progressive formation of ligand holes. This mirrors the systematic evolution of the O K-edge pre-edge observed across the Ni oxide series in Figure 3c. Resonance in the -dominated range in Ni L-edge RPES measurements (Fig. S4c) reveals a spectral attenuation of the ligand-hole hybridised orbitals as cell potential increases. This reduction, highlighted in Fig. S4d (white squares), indicates a progressive electron withdrawal from the ligand-hole states, confirming their direct participation in LNO charge compensation.
The specific nature of these ligand-hole states remains to be directly determined for a comprehensive description of LNO under this formalism. To address this, we now focus on bulk-sensitive Ni L-edge measurements of the pristine LNO electrode, in conjunction with Ni L-edge simulations in the relevant regime. In Figure 4, we start with an electronic ground state assumed to have a fully Ni3+ configuration (3d7) with explicit hybridisation. This case illustrates the influence on the electronic structure of hybridisation alone without spontaneous ligand hole stabilisation. The significant mismatch between the resulting simulated Ni L-edge spectrum and the experimental one further confirms the relevance of the ligand hole formalism over the simple 3d7 model. We continue with considering as the main ligand hole configuration in pristine LNO. When one reads this configuration in the literature, a couple of preconditions are often overlooked. First, the spatial extent of this hybridised orbital is assumed to be confined to a single NiO6 site. And second, due to this single-site spatial extension, the state results from the hybridisation between the Ni 3d eg orbitals and the O ligand orbitals that directly point to the Ni ion as seen in Figure 4a. These conditions lead to a NiO6 layer where ligand configuration is adopted homogeneously by every individual octahedra, as seen in Figure 4b. Figure 4c shows that theoretical calculations based on such a single ligand holes configuration reproduce poorly the measured Ni L3-edge absorption spectrum because these theoretical conditions don’t consider the entire crystallographic structure.
@@
A more realistic approach is considering that the orbital hybridisation expands across the entire structure as in the case of band theory. Indeed, when the Ni -O ligand hybridisation extends across multiple NiO6 octahedra, ligand holes can be more easily stabilised via a “self-regulating response”. [39] This indicates that transferring ligand electrons to the metal is regulated collectively within a set of neighbouring NiO6 octahedra in which the ligand holes reorganise among them via ligand-metal bond alternation, leading to configurations of the type seen in Figure 4e. These configurations have been observed in structural experiments of LNO. [18, 20] As a result, we have ligand configurations of the type , , and that correspond geometrically to unique octahedral configurations. Our hypothesis of the LNO electronic configuration having an equal distribution of these three types replicates well the XAS spectra, as shown in Figure 4f, confirming the self-regulating response is a realistic electronic structure for the pristine LNO electrode. Here, we simulated the Ni L-edge transitions with the three electronic configurations in three independent SIA models. Although the highly covalent yet correlated nature of LNO requires a more complex single Ni L-edge simulation with multiple clusters of NiO6 to properly capture bond-alternation, correlation, and covalency at once, [40, 41] the independent SIA approach herein adopted provides key insights into the basic nature of the Ni L-edge transitions and the corresponding experimental peaks.
@@
Understanding the redox mechanism at the LNO positive electrode side during the charging process ultimately comes down to clarifying how the three , , and ligand configurations adjust during intercalation reactions. In this sense, we have run bulk sensitive Ni L-edge XAS measurements at different cell potentials along the electrochemical profile of half coin cells (also charged at a C-rate of C/20), which are highlighted in Figure 5a. From our previous calculations in Figure 4f we found that the XAS spectra can be interpreted in terms of the intensity of three peaks (at roughly 854, 855 and 856 eV), which denote the LNO crystallographic structures of three different types of NiO6 octahedra and ligand hole configurations. As the battery is charged up to 4.4 V, there is a gradual increase in the intensity of the peak, while the peak is reduced. This suggests that LNO cell charging involves creating new holes that the configuration absorbs. As a result, NiO6 sites reconfigure, Ni-O ligand re-hybridise, and charge redistribute leading to more instances of the , and configurations in alignment with the data in Figure 2. This spectral evolution concentrates in the -dominated Ni L-edge range which in tandem reflects the DMFT-predicted increase of ligand holes, reconciling these fundamentally distinct theoretical pictures used to describe ligand hole evolution. As expected, the discharge process reverses the spectral evolution of the charge process. The filling of these ligand holes dominates the discharge process which we can see in the spectral reduction of the and the increasing spectral features of the configuration after discharging from 4.4 V to 3.0 V vs. LiLi+. A direct consequence of the hole creation/filling dynamics is that electrochemical cycling modulates the distribution of local structures, specifically NiO6 octahedra. We anticipate a coexistence of distinct octahedral geometries associated with the , , and ligand-hole states. A follow-up study will investigate this local structural distribution using operando Ni K-edge XAS to elucidate its dependence on hole formation and (de)intercalation reactions in stoichiometric Ni-rich materials. Instead of counting Ni d electrons, in negative charge transfer compounds, the redox mechanism refers to counting those hole-induced distorted octahedral sites, caused by the dynamical evolution of ligand electronic density.
@@
The precise description of the electronic structure of LMFP64 and LNO has been critical for the distinct charge compensation pathways for these materials. We inferred their electronic structures using the Zaanen-Sawatzky-Allen (ZSA) electronic scheme. It was first discussed by Zaanen, Sawatzky and Allen, [42] where the relative energy positions of the metal d and oxygen p bands was used to categorise correlated transition metal compounds; a more recent work [43] defines four categories based on the material’s relative energy values between and U, which is the on-site Coulomb interaction, i.e., the electron repulsion causing a separation in energy of the 3d-derived electronic states. is defined similarly as in the RPES analysis. The material band gap type formed within the ZSA electronic framework for LMFP64 and LNO dictates the corresponding charge compensation mechanism, illustrating how this band gap can aid in determining the expected material redox nature in general (Figure 6).
@@
Model A in Figure 6 describes the ZSA Mott-Hubbard material class—positive ( U)—applicable to the LMFP64 case. The band gap results from Coulomb repulsion, so it is of a d - d nature. This means that the lowest unoccupied band (conduction band) and highest occupied band (valence band) are both of Fe d character. During cell charge, charge from the Fe 3d valence band will be promoted to the conduction band. The loss of Fe intensity in Figure 1d is a manifestation of this effect, hence, Fe oxidation. Model B in Figure 6 defines the ZSA Negative Charge Transfer material class which determines the LNO electronic structure. Since in these materials, is negative and U positive, the band gap is of a dp - dp type, as shown by our DMFT/RPES analysis in Fig. 2, 3, S2, and S4. During the charge process of LNO cells, electrons are promoted across this dp - dp band gap, in a similar fashion as the spectral migration in Fig. S3, creating holes in the hybridised ligand orbitals while drawing capacity. The LNO system instantaneously undergoes a correlated redistribution of charge resulting in the reconfiguration of the , , and states during electrochemistry. As result, the electrochemical redox of LNO (and related Ni-rich cathodes) can be described by ligand hole formation and annihilation, as shown with Fig. 5.
@@
This ZSA-based redox framework has immediate implications for high-valence Co compounds, which, consistent with recent reports, [44] are expected to exhibit ligand-hole characteristics for Co oxidation states above 3+. In conjunction with the Ni results presented here, this suggests that NMC positive electrodes must display a strongly state-of-charge-dependent ligand-hole contribution. Furthermore, the framework directly informs emerging design rules aimed at stabilising ligand holes in high-valence Fe compounds, again in agreement with recent findings. [45] Beyond these specific chemistries, it establishes a generalised redox strategy applicable to lithium-excess compounds—the systems that originally motivated the study of oxygen redox. Note that the transition from positive to negative charge transfer regimes depends on metal and its formal oxidation state, corresponding to roughly Ni2+, Co3+ and Mn4+. [43] Moreover, the majority of Li-excess compounds are Li-rich Mn-rich. [3] Li2MnO3 and Li2NiO3 reflect the metal and ligand hole endpoints, respectively, consistent with the observed electrochemistry arising from structural and chemical decomposition. [46, 36] Thus, we anticipate these materials could be classed in one of the other two ZSA categories (Positive Charge Transfer or Mixed Valence) with varying p-d band gap types, [16] offering a pathway to study the role of unhybridised oxygen orbitals in delivering reversible capacity without conflicting signals. We acknowledge that structural rearrangements also play a key role in the electrochemical behaviour of Li-rich systems. [22] TM L-edge assisted RPES offers an opportunity to directly explore electronic structures of Li-excess compounds under the ZSA scheme during (de)lithiation and to validate models and descriptions that clarify the role of oxygen oxidation in these compounds.
@@ Conclusion
Here, we present how by experimentally capturing the true nature of the electronic structure of LiMn0.6Fe0.4PO4 and LiNiO2 positive electrode materials under the ZSA scheme, the TM L-edge assisted RPES approach can directly probe their corresponding redox mechanism. Using this framework, we can conclude that the degree of metal-oxygen hybridisation is a key factor in determining the charge compensation mechanism. For more ionic compounds in the positive charge regime, such as the phosphates, electron counting of the d orbital remains valid (e.g. Fe2+/3+ 3d6,3d5). Increasingly covalent systems in the negative charge transfer regime, which includes the LiMO2 (M= Co, Ni) derived family, are better described in terms of electron counting ligand hole states instead (e.g. Ni3+/4+ , ). We expect this framework to also help describe lithium excess compounds. A more accurate probe into the evolution of their electronic structure may resolve the true role of unhybridised oxygen orbitals in delivering reversible capacity in these systems.
Supplementary information The supplementary information includes: Surface versus bulk analysis to identify energy regions in Ni L-edge experiments without surface reconstruction contribution, comparison between Ni L-edge simulations in models with and without explicit orbital hybridisation, integrated experimental and computational data to showcase how Ni oxidation state modulate negative charge transfer energy , and RPES and electrochemical data supporting a ligand hole redox mechanism in LNO.
@@ Acknowledgements
HB acknowledges support from the Royal Society of Chemistry Research Fund Grant R24-8964281967. We acknowledge Diamond Light Source for time on beamline B07 under proposal SI33459-1 and beamline I09 under proposals SI35075-1, SI36917-1, and SI30201. This work was supported by the Faraday Institution’s Degradation (FIRG060, FIRG082) project and the FutureCat (FIRG065, FIRG017) project.
@@
These four powder materials were used during the SI36917 RPES beamtime at the I09 beamline of Diamond Light Source (DLS). Material preparation was performed avoiding air-exposure following the description provided in the RPES section.
@@
The assembly of 2325 coin cells (vs. LiLi+ as a counter electrode) was conducted in the glovebox. A Celgard 2325 Trilayer microporous membrane of 25 m thickness was used as a separator. The cells were filled with 60 L of E151 Solvionic electrolyte, comprised of 12.42:30.82:54.76:2 w/w ratio of LiPF6, ethylene carbonate (EC), ethyl methyl carbonate (EMC), and vinylene carbonate (VC), respectively. To ensure full wetting of the electrodes, the cells were held at rest at 25 ∘C) for 20 hours. Cells underwent one constant-current C/20 (with C = 220 mA/g) charge process to cell potentials (4.0 V, 4.1 V, 4.2 V, 4.3 V, 4.4 V, and 4.6 V) on a Biologic VMP3 potentiostat cycler. The corresponding active mass used to compute specific capacities were 22.15, 23.01, 22.62, 22.5, 22.24, and 22.7 mg.
For each electrochemical condition presented in the manuscript (i.e., each cut-off potential), we assembled and tested three independent coin cells. This approach was adopted to mitigate experimental uncertainties commonly associated with coin cell assembly, particularly electrode misalignment, which can influence overall cell performance. The electrochemical data shown in main text represent the best-performing cell from each set of three. These representative cells were selected based on electrochemical quality (e.g., Coulombic efficiency, voltage profile, and capacity) and alignment with the expected behaviour of LNO at the corresponding state of charge. While variability was observed within some sets, the selected cells follow a consistent trend across the full electrochemical range, reinforcing that they each accurately represent their respective states of charge.
@@
After required cycling coin cells were transferred and carefully disassembled within an argon MBraun glovebox (O 0.1 ppm, H2O 0.1 ppm) to prevent exposure to atmospheric moisture. This guarantees no air-exposure during cell handling after cycling. Once inside the glovebox, we opened the cells to retrieve the electrodes on which later excess electrolyte was thoroughly washed using dimethyl carbonate (DMC) solvent (sourced from Sigma-Aldrich) with an anhydrous purity of 99 percent. After washing and drying, the cleaned electrodes were used in the SI36917 and SI30201 RPES beamline and the SI33459 Ni L-edge beamline after careful sample loading and transportation that avoids air exposure.
@@
LiMn0.6Fe0.4PO4 (LMFP64) powder was purchased from Gelon LIB and used to fabricate electrodes following in-house procedures. A mixture of LMFP64 powder, carbon black, and polyvinylidene fluoride (PVDF) pre-dissolved in N-methyl-2-pyrrolidone (NMP) was prepared in a weight ratio of 93.5:3.5:3.5. Additional NMP was incorporated to achieve the desired solid content. The resulting slurry was then applied to carbon-coated aluminium foil, dried overnight at 120 ∘C, and calendared to a final density of 2.15 g cm-3. To charge the material to a specific state of charge (SoC), half coin cells were assembled with LMFP64 positive electrodes (14.8 mm diameter), a polymer separator (H2325), LP57 electrolyte, and a lithium metal disc as the counter electrode. The cells were then charged galvanostatically to a specified potential using a current of 7.9307 mA/g (equivalent to a C/20 charge rate with and average of 22 mg of active material). Specific capacity reported in the manuscript refer to the active material mass at 21.6, 21.8, 21.5, 22.1, 21.5, and 21.7 mg for the 0%, 20%, 40%, 70%, and 100% SoC respectively.
Similar to the LNO case, we assembled and tested three independent coin cells for each electrochemical condition presented in the manuscript (i.e., each cut-off potential). This approach mitigates experimental uncertainties commonly associated with coin cell assembly. The electrochemical data shown in main text represent the best-performing cell from each set of three. These representative cells were selected based on electrochemical quality (e.g., Coulombic efficiency, voltage profile, and capacity) and alignment with the expected behaviour of LMFP64 at the corresponding state of charge. The selected cells follow a consistent trend across the full electrochemical range, reinforcing that they each accurately represent their respective states of charge.
After charging, the cells were carefully transferred and disassembled in an Ar-filled glovebox (O 0.1 ppm, H2O 0.1 ppm), where they were rinsed with DMC and dried. This material was used in the SI35075 RPES beamtime after careful sample loading and transportation that avoids air exposure.
@@
Resonant Photoemission Spectroscopy (RPES) at Ni 2p and Fe 2p absorption thresholds were recorded at beamline I09 DLS across three sessions: Ni RPES in SI30201 and SI36917 beamtimes, and Fe RPES in SI35075 beamtime. For all these sessions, we used a VG Scienta EW4000 detector with a 70 frame/sec CCD camera. The four Ni-based materials in the formed of powder were used in this beamline. LNO half-cells, described in the LNO coin cell electrochemical cycling section, charged to and opened at 4.2 V, 4.4 V, and 4.6 V were used in these beamline. Additionally, LMFP electrodes, described in the LMFP cell section and charged at 20%, 40%, 70%, and 100% SoC were also used in this beamline. The total-energy resolution for the measurements at I09 was 0.2 eV. The calibration of the photon energy was performed by comparing the kinetic energy of the Au 4f peak. The base pressure during all measurements was less than 210-10 torr and all measurements were performed at room temperature. After the LNO and LMFP cells were transferred into the argon MBraun glovebox (O 0.1 ppm, H2O 0.1 ppm) and excess of electrolyte were washed off with DMC, we prepared the material for the RPES beamtimes inside the WMG glovebox. We mounted the samples on copper plates and placed these plates into the sample cassette provided by the I09 beamline team. This cassette features a sealing system that preserves the argon-inert environment during transport and loading at the I09 facility. It ensures the samples remain protected from environmental contamination and air exposure throughout the process, from the electrochemical cycling to introduction into the spectrometer sampling chamber.
@@
Ni L-edge X-ray Absorption Spectroscopy were measured in TEY and total fluorescent yield (TFY) mode at the B07 beamline [51] at DLS under the SI33459 beamtime. Energy calibration was performed using a NiO reference. The Ni spectra were collected with a energy resolution 0.15. LNO half-cells, described in the LNO coin cell electrochemical cycling section, charged to and opened at 4.0 V, 4.1 V, 4.2 V, 4.3 V, and 4.4 V were used in these beamline. These electrodes were retrieved, DMC-cleaned, and prepared under Ar-inert conditions in our Ar-filled glovebox (O 0.1 ppm, H2O 0.1 ppm) at WMG. We proceeded to load the samples onto sample plates provided by the B07 team. Inside the glovebox, the plates were vacuum-sealed in pouch cells designed to hold the Ar-environment long enough to allow for a secure transportation and loading once at the B07 facilities, guaranteeing no-air exposure during the operation and transport of the material.